OpenStudy (anonymous):

0! = 0 True or False?

OpenStudy (amistre64):

well, we are told by definition what 0! is

OpenStudy (anonymous):

True

OpenStudy (anonymous):

false

OpenStudy (anonymous):

i promise it is false 0!=1

OpenStudy (antonio_xx2):

false

OpenStudy (anonymous):

Would false be correct?

OpenStudy (anonymous):

yes false is the right answer

OpenStudy (anonymous):

Yes my bad

OpenStudy (anonymous):

False is correct

OpenStudy (anonymous):

False Already .

OpenStudy (anonymous):

0!= 1 would be .

OpenStudy (accessdenied):

Most consider the zero factorial to be like an empty product. 0! = 1 a bit like how a^0 = 1 (for a not 0). There are no factors present, so the only remaining item is the multiplicative identity 1. Another way of viewing it might be 1! = 1* 0!, or 1 = 1 * 0! = 0!. Just giving some ideas why rather than a true/false answer. :)